Several years ago, Geneva purchased an office building for $1,000,000. She sold the building this year for $1,250,000. Accumulated depreciation on the building totaled $965,000 at the date of sale. Assume Geneva’s total income is such that her long-term capital gains rate is 20%. How much tax will she owe in connection with sale of the building?
Answer #1

Ans $ 243000

Profit on Sale = Sales Value - Written down value

= 1250000 - ( purchase cost - accumulated depreciation)

= 1250000 - (1000000 - 965000)

= 1215000

Tax on Profit = 1215000 * 20%

= $ 243000